OpenStudy (anonymous):

Media Services charges $20 for a phone and $30/month for its economy plan. The equation c=30t+20 describes the total cost, c, of operating a Media Services phone for phone for t months. The total cost for 2 months of service is $ If a customer has only $90 available, how many months of service can she receive?

OpenStudy (anonymous):

the total cost for two months is if c=30(t)+20 ,t is the variable for months therefore sub t=2 in eqn c=30(2)+20 =80 to find months of service, make t the subject t=c-20/30 t= (90 - 20 )/30 t = 2 1/3